A= ) = 0 S NEN (b) B= (d) Da LEE O NO

Obtain the inverse of each matrix.

part a only
Answer #1

Ans) Inverse of matrix

IQ cowo Olow! rowe orao AT= o7 Crow becomes on T 1 column & Column becomes - Row) 7 find determinant of each of axa winos mat

= Determinant of Original Matrix A = 3(-3-10) - 062-0)+ 4(1940) = 3x -13 - 0 + 40 = -39 +40 = 1 => I A-= Adj (A) I x det (A)